A laboratory employee is mixing a 10% saline solution with 4% saline solution. How much of each solution is needed to make 500 milliliters of a 6% solution.
Answer by mananth(16946)

A laboratory employee is mixing a 10% saline solution with 4% saline solution. How much of each solution is needed to make 500 milliliters of a 6% solution.
percent ---------------- quantity
10 ---------------- x
4 ---------------- 500-x
6.00% ---------------- 500
10x+4(500-x)=6*500
10x+2000 -4x=3000
10x-4x=3000-2000
6x=1000
/6
x=166.67ml 10% solution
Balance 4% = 333.33 ml